OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term?

OpenStudy (anonymous):

multiply eq P by 2 and subtract from eq Q. x will got eliminated.

OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term? 2(x + y = 8) −2(x + y = 8) −1(2x + 5y = 24) −2(2x + 5y = 24)

OpenStudy (anonymous):

A ?

OpenStudy (anonymous):